Milton Company uses a standard cost system.

  • Standards call for 2 direct labor hours per widget. The standard labor rate is $10 per hour.
  • Four pounds of material are needed for each widget at a standard cost of $2.25 per pound.

Milton actually produces 17,000 widgets during the year.

  • Direct labor costs are $359,100 for 34,200 hours.
  • Milton purchased and used 70,000 pounds of materials at $2.50 per pound.

For direct labor, what is the rate (price) variance? Answers are labeled as favorable (F) or unfavorable (U).

A.

$34,200 U

B.

$17,100 U

C.

$8,500 U

D.

$19,100 U

E.

$17,000 F
Answer #1

Labor rate variance

= (SR-AR) * AH

= (10 - 359,100/34,200) * 34,200

= (10 - 10.5) * 34,200

= 17,100 U

OPTION B